doubledayemilie:

Is (x − 2) a factor of f(x) = x3 − 2x2 + 2x + 3? Use either the remainder theorem or the factor theorem to explain your reasoning.

doubledayemilie:

do you mind helping me ?

sterling:

what grade go to mathway it would help

doubledayemilie:

11 i tried that but i didnt work

sterling:

microsoft math solver

doubledayemilie:

tried that too

doubledayemilie:

that has different numbers its not the same numbers

hiiiii2221:

did u try brainly

doubledayemilie:

yeah still has the different numbers

hiiiii2221:

just substitute it than

hiiiii2221:

change out the numbers but do the same thing as they did to get their answer

doubledayemilie:

okay thank you

surjithayer:

put x-2=0 x=2 \[f(2)=2^3-2(2)^2+2 \times 2+3\] \[f(2)=7 \neq 0\] Hence by remainder theorem x-2 is not a factor of f(x)